# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v1

@@ III-A Cluster Memory Contrastive Baseline
To facilitate the description of our method, we first introduce the notation. Let $X_{d}=\{x^{d}_{i}\}_{i=1}^{N}$ denote a set of drone images with $N$ instances, and $X_{s}=\{x^{s}_{i}\}_{i=1}^{M}$ denote a set of satellite images with $M$ instances. $F_{d}=\{\mathbf{f}^{d}_{i}\}_{i=1}^{N}$ and $F_{s}=\{\mathbf{f}^{s}_{i}\}_{i=1}^{M}$ represent the feature representations extracted by the shared feature extractor $\mathcal{F}_{backbone}$ for the drone and satellite views, respectively. $q^{d}$ and $q^{s}$ denote the drone query instance features and satellite query instance features extracted by the feature extractor $\mathcal{F}_{backbone}$. The processing procedure of the backbone can be described as follows.
@@
Cross-View Feature Memory Initialization. At the beginning of each training epoch, the feature memories for both views are initialized. Specifically, the cluster representations of the drone view $\{\phi_{1}^{d},\phi_{2}^{d},\cdots,\phi_{K}^{d}\}$ and the satellite view $\{\phi_{1}^{s},\phi_{2}^{s},\cdots,\phi_{L}^{s}\}$ are stored in the drone memory dictionary $\mathcal{M}_{d}$ and satellite memory dictionary $\mathcal{M}_{s}$, respectively, for view-specific memory initialization. This process can be formalized as:
$\mathcal{M}_{d}\ \leftarrow\{\phi_{1}^{d},\phi_{2}^{d},\cdots,\phi_{K}^{d}\}$ (2)
$\mathcal{M}_{s}\ \leftarrow\{\phi_{1}^{s},\phi_{2}^{s},\cdots,\phi_{L}^{s}\}$ (3)
$\phi_{k}^{d}=\frac{1}{|\mathcal{H}_{k}^{d}|}\sum_{\mathbf{f}_{n}^{d}\in\mathcal{H}_{k}^{d}}\mathbf{f}_{n}^{d}$ (4)
$\phi_{l}^{s}=\frac{1}{|\mathcal{H}_{l}^{s}|}\sum_{\mathbf{f}_{m}^{s}\in\mathcal{H}_{l}^{s}}\mathbf{f}_{m}^{s}$ (5)
where, $n=1,\cdots,N$, $m=1,\cdots,M$, $k=1,\cdots,K$, $l=1,\cdots,L$, $\mathcal{H}_{k(l)}^{d(s)}$ denotes the $k$-th cluster set in drone or satellite view, ||$\cdot$|| ⋅ | indicates the number of instances per cluster.
Cross-View Feature Memory Updating. During the training phase, we randomly sample $P$ distinct localization points from the training set, with each point comprising a fixed number $Z$ of instances. This sampling strategy results in a minibatch containing a total of $P\times Z$ query images. Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism. Through this approach, we iteratively update the features for both perspectives, as formalized in Eq.6 and Eq.7.
@@
where, $\tau$ denotes the temperature hyper-parameter. The cluster centroids $\phi^{d(s)}_{k(l)}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance $q_{d(s)}$ and all clusters. Here, $\phi^{d(s)}_{+}$ corresponds to the positive cluster feature associated with $q_{d(s)}$.
@@ III-B Dynamic Hierarchical Memory Learning
In CVGL tasks, baseline networks can rapidly respond to local environmental changes and learn subtle differences in recent scenes. However, they are susceptible to transient noise and local anomalies, leading to instability in feature representation. Moreover, due to the complexity introduced by geographic variations, baseline networks struggle to extract globally consistent and stable features effectively. To address these challenges, we propose a Dynamic Hierarchical Memory Learning (DHML) method to achieve a balance between instantaneous adaptation and long-term stability.
Consistent with the baseline network, we construct the dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example. To build the long-term memory, we first construct two novel clustering memory dictionaries according to Eq.2 and Eq.3, which are used to store the short-term $\mathcal{M}_{d}^{s}$ and long-term $\mathcal{M}_{d}^{l}$ memories, respectively.
@@
Unlike the weight updating scheme in Eq.6 and Eq.7, we assign different weights to the input features and the historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, during the initialization of the long-term memory update we deliberately reduce the influence of the current input (i.e., the input weight is $1-\alpha$) and amplify the influence of the historical memory (i.e., the historical weight is $\alpha$) to maintain consistency and stability of the memory. The update procedure can be defined as follows:
@@
In order to learn local input information while preserving global consistency, we further introduce an adaptive weight updating mechanism to update the long-term memory. In detail, we first compute the Euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range $[0,1]$ via the sigmoid function, yielding an adaptive update coefficient $\beta$:
@@
This adaptive updating strategy ensures that, when there is a significant deviation between the input and the short-term memory, the current input information exerts a greater influence on the long-term memory, thereby enabling a more rapid capture of scene changes. Conversely, when the deviation is small, the historical memory predominantly contributes to maintaining global consistency. Finally, the long-term and short-term memories are fused in a weighted manner to yield the combined feature $\varphi_{k}^{db}$ representation:
@@
where $w_{l}$ and $w_{s}$ denote the respective fusion weights. As a result, we obtain a comprehensive memory dictionary $\mathcal{M}_{d}^{cb}$ that effectively integrates long-term global consistency with localized short-term adaptability, ensuring robust and stable feature representation across varying environmental conditions.
@@
where, $\tau$ denotes the temperature hyper-parameter. The combined cluster centroids $\phi^{db}_{k}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance $q^{d}$ and all clusters. Here, $\phi^{db}_{+}$ corresponds to the positive cluster feature associated with $q^{d}$.
To further reinforce local adaptability while maintaining stability, we integrate our proposed loss function with the baseline loss $\mathcal{L}^{d}_{cv}$, ensuring that the model effectively learns fine-grained scene variations while preserving robust global consistency. The final optimization objective is formulated as:
$\mathcal{L}_{dhml}^{d}=\mathcal{L}^{d}_{cv}+\lambda_{cv}\mathcal{L}^{dd}_{cv}$ (18)
where $\lambda_{cv}$ is a trade-off parameter that controls the contribution of the $\mathcal{L}^{d}_{dhml}$.The satellite contrastive loss $\mathcal{L}_{dhml}^{s}$ are obtained by the same way.
We effectively refine the feature representation by balancing cross view local adaptability and global stability by combining cross view hierarchical memory learning approach:
@@ III-C Information Consistency Evolution Learning
By integrating the baseline network with the dynamic hierarchical memory learning module, the proposed framework exhibits a certain capability for cross-view consistency learning. However, both the baseline network and the dynamic hierarchical memory learning module primarily focus on intra-view feature learning, even though the drone-view and satellite-view features share a common backbone network. As a result, the deep correlations between features from different views remain insufficiently explored, limiting the model’s ability to fully mitigate and bridge the challenges posed by drastic viewpoint variations.
To address this limitation, we propose an Information Consistency Evolution Learning (ICEL) module, as illustrated in Fig.3. This module leverages a threshold-based filtering and ranking-driven neighborhood selection strategy to reinforce intra-view consistency while simultaneously enabling cross-view feature alignment. By facilitating consistency evolution across views, ICEL enhances the discriminative power and relational modeling of both intra-view and cross-view feature representations, thereby improving the model’s robustness in complex geo-localization scenarios.
In the ICEL module, unlike the baseline and DHML module, we employ instance-level features instead of cluster-level representations. To achieve this, we define two instance memory dictionaries, $\mathcal{I}_{d}$ and $\mathcal{I}_{s}$, which store the instance features extracted by the shared feature encoder, as described in Section III-A. This process can be formally expressed as follows:
@@
In addition, we define the similarity between a given drone (or satellite) query $q_{z}^{d{(s)}}$ and each drone (or satellite) instance $\mathbf{f}_{v}^{d{(s)}}$ in the training set as:
@@
Based on the instance memory dictionary and similarity computation (as shown in Fig.3), we adopt two learning modes: intra-view representation enhancement and cross-view feature association. To ensure that the learning process relies only on high-confidence neighborhood information, we introduce a neighborhood selection strategy that integrates threshold-based filtering and top-$k$ ranking.he threshold-based filtering strategy can be defined as follows:
@@
where, $\Omega^{dd}$ represents the set of drone feature instances selected as valid intra-view neighbors and $\Omega^{ds}$ denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features. $\gamma\in(0,1)$ is a predefined threshold. The $\Omega^{ss}$ and $\Omega^{sd}$ can be calculated by the same way.
Although fixed-threshold-based neighborhood selection identifies highly relevant samples, it may still yield an excessively large neighborhood due to its global filtering nature. To refine this into a more compact and discriminative set, we introduce a top-$k$ ranking selection strategy. Specifically, we independently select the top-$k_{1}$ most relevant samples to construct a strictly filtered neighborhood, ensuring high-confidence associations, while also selecting the top-$k_{2}$ samples define an expanded neighborhood, capturing additional informative relationships. The process can be defined by:
@@
Based on the constructed neighborhoods, we further define three loss functions. First, within a larger-scale neighborhood, to enhance the aggregation of similar samples in both intra-view and cross-view settings, we compute the normalized similarity distribution over the selected neighborhood set and optimize it using a cross-entropy loss to maximize the alignment between the query sample and its neighborhood samples. The intra-view $\mathcal{L}_{\Omega}^{dd}$ and cross-view $\mathcal{L}_{\Omega}^{ds}$ losses can be formulated as:
@@
Incorporating neighborhoods $\mathcal{N}_{k1}^{dd}$, $\mathcal{N}_{k2}^{dd}$ and $\mathcal{N}_{k1}^{ds}$, $\mathcal{N}_{k2}^{ds}$, in addition to intra-view and cross-view neighborhood contrastive losses, we introduce consistency loss and mutual information loss to enhance feature robustness and semantic alignment across views.
The consistency loss minimizes the Kullback-Leibler to enforce alignment of relative similarity distributions within the large-scale neighborhood $k_{2}$ across both intra-view and cross-view representations. This ensures that features from different modalities or viewpoints maintain a consistent neighborhood structure. The intra-view loss $\mathcal{L}_{k2}^{dd}$ and cross-view loss $\mathcal{L}_{k2}^{ds}$ under this consistency loss can be formulated as:
$p_{z,v}^{dd}=\frac{\exp(S(q_{z}^{d},\mathbf{f}_{v}^{d}))}{\sum_{v^{\prime}\in\mathcal{N}_{k2}^{d}}\exp(S(q_{z}^{d},\mathbf{f}_{v^{\prime}}^{d}))}$ (31)
@@
where, $p_{z,v}^{d(s)}$ is the softmax-normalized similarity distribution, $v^{\prime}$ is used for summation to normalize the similarity scores over all neighborhood samples. $\mathcal{L}_{k2}^{ss}$ and $\mathcal{L}_{k2}^{sd}$ can be obtained in the same way.
Finally, within a more restrictive and smaller set of $k_{1}$ neighbors, a mutual information constraint is proposed to learn fine-grained relationships between the query feature and local samples. This complements the consistency constraint described above, achieving a balance between global structural alignment and local discriminative information. Within the selected $k_{1}$ neighborhood, we first compute, for each neighbor sample $\mathbf{f}_{v_{i}}^{d(s)}$ (where $i=1,\ldots,k_{1}$), the similarity between the deep query feature $q_{z}^{d}$ and the neighbor feature $\mathbf{f}_{v_{i}}^{d(s)}$ using the similarity function $S(q_{z}^{d(s)},\mathbf{f}_{v_{i}}^{d(s)})$. Based on these similarity scores, the posterior probability for each neighbor is defined via a softmax normalization as:
@@
where $p_{i}^{dd}(v_{i}|q_{z}^{d})$ and $p_{i}^{ds}(v_{i}|q_{z}^{d})$ represent the normalized similarity distributions over the selected $R$ intra-view and cross-view neighbors, respectively. $\mathcal{L}_{k1}^{ss}$ and $\mathcal{L}_{k1}^{ss}$ can be obtain in the same way. These mutual information losses maximize the entropy of the learned feature distributions, ensuring that the query features maintain a well-separated and informative representation within their respective local neighborhoods.
For each viewpoint, the corresponding loss can be formulated as follows:
$\mathcal{L}_{icel}^{dd}=\mathcal{L}_{\Omega}^{dd}+\lambda_{k1}\mathcal{L}_{k1}^{dd}+\lambda_{k2}\mathcal{L}_{k2}^{dd}$ (38)
$\mathcal{L}_{icel}^{ds}=\mathcal{L}_{\Omega}^{ds}+\lambda_{k1}\mathcal{L}_{k1}^{ds}+\lambda_{k2}\mathcal{L}_{k2}^{ds}$ (39)
$\mathcal{L}_{icel}^{ss}=\mathcal{L}_{\Omega}^{ss}+\lambda_{k1}\mathcal{L}_{k1}^{ss}+\lambda_{k2}\mathcal{L}_{k2}^{ss}$ (40)
$\mathcal{L}_{icel}^{sd}=\mathcal{L}_{\Omega}^{sd}+\lambda_{k1}\mathcal{L}_{k1}^{sd}+\lambda_{k2}\mathcal{L}_{k2}^{sd}$ (41)
where $\lambda_{k1}$ and $\lambda_{k2}$ are importance adjustment factors. The final optimization for neighbor learning is denoted by the following combination:
$\mathcal{L}_{icel}=\mathcal{L}_{icel}^{d}+\mathcal{L}_{icel}^{s}$ (42)
where $\mathcal{L}_{icel}^{d}=\mathcal{L}_{icel}^{dd}+\mathcal{L}_{icel}^{ds}$ and $\mathcal{L}_{icel}^{s}=\mathcal{L}_{icel}^{ss}+\mathcal{L}_{icel}^{sd}$. The overall loss for DMNIL is denoted as:

### v2

@@ III-A Dual-Path Contrastive Learning Baseline
We first introduce the notations used in this paper for better clarity. Let $X_{d}=\{x^{d}_{i}\}_{i=1}^{N}$ denote a set of drone images with $N$ instances, and $X_{s}=\{x^{s}_{i}\}_{i=1}^{M}$ denote a set of satellite images with $M$ instances. $F_{d}=\{\mathbf{f}^{d}_{i}\}_{i=1}^{N}$ and $F_{s}=\{\mathbf{f}^{s}_{i}\}_{i=1}^{M}$ represent the feature representations extracted by the shared feature extractor $\mathcal{F}_{backbone}$ for the drone and satellite views, respectively. $q^{d}$ and $q^{s}$ denote the drone query instance features and satellite query instance features extracted by the feature extractor $\mathcal{F}_{backbone}$. The processing procedure of the backbone can be described as
@@
Cross-View Feature Memory Initialization. At the beginning of each training epoch, the feature memories for both views are initialized. Specifically, the cluster representations of the drone view $\{\phi_{1}^{d},\phi_{2}^{d},\cdots,\phi_{K}^{d}\}$ and the satellite view $\{\phi_{1}^{s},\phi_{2}^{s},\cdots,\phi_{L}^{s}\}$ are stored in the drone memory dictionary $\mathcal{M}_{d}$ and satellite memory dictionary $\mathcal{M}_{s}$, respectively, for view-specific memory initialization. This process can be formalized as
$\mathcal{M}_{d}\ \leftarrow\{\phi_{1}^{d},\phi_{2}^{d},\cdots,\phi_{K}^{d}\},\mathcal{M}_{s}\ \leftarrow\{\phi_{1}^{s},\phi_{2}^{s},\cdots,\phi_{L}^{s}\}$ (2)
$\phi_{k}^{d}=\frac{1}{|\mathcal{H}_{k}^{d}|}\sum_{\mathbf{f}_{n}^{d}\in\mathcal{H}_{k}^{d}}\mathbf{f}_{n}^{d},\phi_{l}^{s}=\frac{1}{|\mathcal{H}_{l}^{s}|}\sum_{\mathbf{f}_{m}^{s}\in\mathcal{H}_{l}^{s}}\mathbf{f}_{m}^{s}$ (3)
where $n=1,\cdots,N$, $m=1,\cdots,M$, $k=1,\cdots,K$, $l=1,\cdots,L$, $\mathcal{H}_{k(l)}^{d(s)}$ denotes the $k$-th cluster set in drone or satellite view, ||$\cdot$|| ⋅ | indicates the number of instances per cluster.
Cross-View Feature Memory Updating. During the training phase, we randomly sample $P$ distinct localization points from the training set, with each point comprising a fixed number $Z$ of instances. This sampling strategy results in a minibatch containing a total of $P\times Z$ query images. Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism. Through this method, we iteratively update the features for both perspectives. This process can be formalized as
@@
where $\tau$ denotes the temperature hyper-parameter. The cluster centroids $\phi^{d(s)}_{k(l)}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance $q_{d(s)}$ and all clusters. Here, $\phi^{d(s)}_{+}$ corresponds to the positive cluster feature associated with $q_{d(s)}$.
@@ III-B Dynamic Hierarchical Memory Learning
Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions display significant differences. However, such spatial characteristics often introduce ambiguity in intra-view feature representations. To address this, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
Consistent with the baseline network, we construct the dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example. To build the long-term memory, we first construct two novel clustering memory dictionaries according to Eq.2, which are used to store the short-term $\mathcal{M}_{d}^{s}$ and long-term $\mathcal{M}_{d}^{l}$ memories, respectively.
@@
Unlike the weight updating scheme in Eq.4 and Eq.5, we assign different weights to the input features and the historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates. The update procedure can defined as
@@
Additional, we further introduce an adaptive weight updating mechanism to update the long-term memory. In detail, we compute the Euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range $[0,1]$ via the sigmoid function, yielding an adaptive update coefficient $\beta$:
@@
Then, the long-term and short-term memories are fused in a weighted manner to yield the combined feature $\varphi_{k}^{db}$ representation:
@@
where $w_{l}$ and $w_{s}$ denote the respective fusion weights. As a result, we obtain a comprehensive memory dictionary $\mathcal{M}_{d}^{cb}$ that effectively integrates long-term global consistency with localized short-term adaptability.
@@
where $\tau$ denotes the temperature hyper-parameter. The combined cluster centroids $\phi^{db}_{k}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance $q^{d}$ and all clusters. Here, $\phi^{db}_{+}$ corresponds to the positive cluster feature associated with $q^{d}$.
Finally, we introduce a fusion factor $\lambda_{cv}$ to perform weighted fusion of $\mathcal{L}^{dd}_{id}$ and $\mathcal{L}^{d}_{cv}$, aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss $\mathcal{L}_{dhml}^{d}$. Similarly, $\mathcal{L}_{dhml}^{s}$ can be obtained in the same way. Therefore, the dynamic hierarchical memory learning process can be summarized as
@@ III-C Information Consistency Evolution Learning
The baseline network and the DHML module primarily focus on intra-view feature learning, lacking the ability to model and associate cross-view features effectively. Therefore, the proposed ICEL module employs a threshold-based and ranking-based neighborhood selection strategy to achieve effective cross-view feature alignment. By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and representation capability of cross-view features.
In the ICEL module, unlike the baseline and DHML module, we employ instance-level features instead of cluster-level representations. To achieve this, we define two instance memory dictionaries, $\mathcal{I}_{d}$ and $\mathcal{I}_{s}$, which store the instance features extracted by the shared feature encoder, as described in Section III-A. This process can be described as
@@
In addition, we define the similarity between a given drone (or satellite) query $q_{z}^{d{(s)}}$ and each drone (or satellite) instance $\mathbf{f}_{v}^{d{(s)}}$ in the training set as
@@
Then, we adopt a threshold-based neighborhood selection strategy to select reliable neighbors, which can be defined as
@@
where $\Omega^{dd}$ represents the set of drone feature instances selected as valid intra-view neighbors. $\Omega^{ds}$ denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features. $\gamma\in(0,1)$ is a predefined threshold. The $\Omega^{ss}$ and $\Omega^{sd}$ can be obtained by the same way.
Additionally, we introduce a ranking-based neighborhood selection strategy. Specifically, we independently select the top-$k_{1}$ most relevant samples to construct a strictly filtered neighborhood, ensuring high-confidence associations, while also selecting the top-$k_{2}$ samples define an expanded neighborhood, capturing additional informative relationships. The process can be defined as
@@
Based on the constructed neighborhoods, the normalized similarity distribution is computed within the larger neighborhood to promote the aggregation of similar samples in both intra-view and cross-view scenarios. This distribution is then optimized by a cross-entropy loss to enhance alignment between query samples and their neighbors. The intra-view $\mathcal{L}_{\Omega}^{dd}$ and cross-view $\mathcal{L}_{\Omega}^{ds}$ losses can be formulated as
@@
Incorporating neighborhoods $\mathcal{N}_{k1}^{dd}$, $\mathcal{N}_{k2}^{dd}$ and $\mathcal{N}_{k1}^{ds}$, $\mathcal{N}_{k2}^{ds}$, we introduce consistency loss to enhance alignment of relative similarity distributions within the large-scale neighborhood $k_{2}$ across both intra-view and cross-view representations. The intra-view loss $\mathcal{L}_{k2}^{dd}$ and cross-view loss $\mathcal{L}_{k2}^{ds}$ under this consistency loss can be formulated as
$p_{z,v}^{dd}=\frac{\exp(S(q_{z}^{d},\mathbf{f}_{v}^{d}))}{\sum_{v^{\prime}\in\mathcal{N}_{k2}^{dd}}\exp(S(q_{z}^{d},\mathbf{f}_{v^{\prime}}^{d}))}$ (28)
@@
where $p_{z,v}^{d(s)}$ is the softmax-normalized similarity distribution, $v^{\prime}$ is used for summation to normalize the similarity scores over all neighborhood samples. $\mathcal{L}_{k2}^{ss}$ and $\mathcal{L}_{k2}^{sd}$ can be obtained in the same way.
Within a more restrictive and smaller set of $k_{1}$ neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above, achieving a balance between global structural alignment and local discriminative information. Within the selected $k_{1}$ neighborhood, we first compute, for each neighbor sample $\mathbf{f}_{v_{i}}^{d(s)}$ (where $i=1,\ldots,k_{1}$), the similarity between the deep query feature $q_{z}^{d}$ and the neighbor feature $\mathbf{f}_{v_{i}}^{d(s)}$ using the similarity function $S(q_{z}^{d(s)},\mathbf{f}_{v_{i}}^{d(s)})$. Based on these similarity scores, the posterior probability for each neighbor is defined via a softmax normalization as:
@@
where $p_{i}^{dd}(v_{i}|q_{z}^{d})$ and $p_{i}^{ds}(v_{i}|q_{z}^{d})$ represent the normalized similarity distributions over the selected $R$ intra-view and cross-view neighbors, respectively. $\mathcal{L}_{k1}^{ss}$ and $\mathcal{L}_{k1}^{ss}$ can be obtain in the same way.
For each view combination, the loss is computed as the weighted sum of three components: the original neighborhood loss $\mathcal{L}_{\Omega}$, the strict neighborhood loss $\mathcal{L}_{k1}$, and the extended neighborhood loss $\mathcal{L}_{k2}$:
$\mathcal{L}_{icel}^{VQ}=\mathcal{L}_{\Omega}^{VQ}+\lambda_{k1}\mathcal{L}_{k1}^{VQ}+\lambda_{k2}\mathcal{L}_{k2}^{VQ},\quad V,Q\in\{d,s\}$ (35)
where $V$ and $Q$ denote the drone or satellite view.
The total neighbor learning loss is then formulated as
$\mathcal{L}_{icel}=\sum_{V\in\{d,s\}}\sum_{Q\in\{d,s\}}\mathcal{L}_{icel}^{VQ}$ (36)
Finally, the overall loss of our method is defined as